# Pauli Propagation for Imaginary Time Evolution

### v1

@@ Abstract
We extend the Pauli Propagation framework to simulate imaginary time evolution. By deriving explicit update rules for the propagation of Pauli operators under imaginary time evolution generated by Pauli strings, we introduce an imaginary time Pauli Propagation (ITPP) algorithm for approximating imaginary time dynamics directly in the Pauli basis. This approach enables the computation of thermal and ground-state properties while retaining the key computational advantages of Pauli Propagation. Benchmarking ITPP on the one-dimensional transverse-field Ising model demonstrates that truncation provides a controlled trade-off between accuracy and computational cost, while also revealing challenges associated with operator growth under imaginary time evolution. Finally, combining imaginary time and real-time Pauli Propagation naturally suggests a pathway toward simulating open quantum system dynamics within a unified framework.
@@ I Introduction
Classical simulation of quantum many-body systems is notoriously challenging due to the exponential growth of the Hilbert-space dimension with system size, rendering exact diagonalization rapidly intractable. As a consequence, approximate numerical methods are indispensable. Established techniques such as tensor network algorithms [Schollw_ck_2011, Or_s_2014] and quantum Monte Carlo [RevModPhys.73.33, Assaad2008] have enabled substantial progress, yet they also face well-known limitations, including entanglement growth in out-of-equilibrium dynamics [RevModPhys.82.277, Calabrese_2005] and sign-problem constraints in frustrated or fermionic models [PhysRevB.41.9301, PhysRevLett.94.170201].
@@
The original formulation of Pauli Propagation applies to real-time dynamics. In this work, we extend Pauli Propagation to imaginary time evolution. Imaginary time evolution (ITE) is a fundamental tool for state preparation, enabling access to thermal states and filtering toward ground states, and is routinely implemented in classical simulation methods such as quantum Monte Carlo [RevModPhys.73.33] and tensor-network algorithms [Schollw_ck_2011]. We derive explicit rules governing the evolution of Pauli operators under imaginary time evolution generated by arbitrary Pauli strings. Building on these rules, we construct an algorithm for approximating thermal and ground-state observables. As in real-time Pauli Propagation, the number of Pauli terms typically grows rapidly under imaginary time evolution, requiring truncation strategies to retain computational feasibility. Moreover, the ability to handle non-unitary imaginary time evolution within the Pauli Propagation framework opens the door to combining imaginary- and real-time dynamics, thus enabling the simulation of open quantum system dynamics [PRXQuantum.3.010320].
To benchmark the method and to assess its practical performance, we apply the imaginary time evolution Pauli Propagation algorithm (ITPP) to the one-dimensional transverse-field Ising model in the ordered phase. We compare its results with Trotterized imaginary time evolution and exact calculations, and we examine several truncation strategies, including coefficient thresholds and fixed-rank schemes.
The paper is organized as follows. In Sec. II, we briefly review imaginary time evolution, thermal states, and the standard Pauli Propagation algorithm for real-time dynamics. Next, we introduce our extension of Pauli Propagation to imaginary time dynamics. Numerical experiments are presented in Sec. III. Finally, Sec. IV summarizes our findings and discusses possible future directions.
@@ II Theory
In this section, we review the theoretical background underlying our approach. We begin by summarizing imaginary time evolution and its connection to thermal and ground states. We then outline the standard Pauli Propagation algorithm for real-time dynamics. Finally, we introduce our extension of Pauli Propagation to imaginary time evolution, which forms the basis of the method used in this work.
II.1 Imaginary Time Evolution
@@
More generally, imaginary time evolution can be formulated at the level of density matrices. Given an initial density matrix $\rho_{0}$, its imaginary time evolved state is defined as
$\rho(\tau)=\frac{e^{-\tau\hat{H}}\,\rho_{0}\,e^{-\tau\hat{H}}}{\operatorname{Tr}\!\left[e^{-2\tau\hat{H}}\rho_{0}\right]}.$ (2)
This expression ensures proper normalization, $\operatorname{Tr}[\rho(\tau)]=1$, at all imaginary times.
@@
In this case, the imaginary time evolution defined in Eq. (2) yields, for an arbitrary imaginary time $2\tau$, the thermal state
$\rho_{\mathrm{th}}(\tau)=\frac{e^{-2\tau\hat{H}}}{\operatorname{Tr}\!\left[e^{-2\tau\hat{H}}\right]}$, (4)
obtained by evolving the infinite-temperature state at $\tau=0$ to imaginary time $2\tau$, which sets the inverse temperature. In the zero-temperature limit, $\tau\to\infty$, the imaginary time evolution converges to a projector onto the ground state.
Expectation values of observables $\hat{O}$ at imaginary time $2\tau$ are obtained as
$\langle\hat{O}\rangle_{\tau}=\operatorname{Tr}\!\left[\hat{O}\,\rho(\tau)\right]=\frac{\operatorname{Tr}\!\left[\hat{O}\,e^{-\tau\hat{H}}\right]}{\operatorname{Tr}\!\left[e^{-\tau\hat{H}}\right]}$. (5)
In the zero-temperature limit $\tau\to\infty$, the thermal state $\rho(\tau)$ projects onto the ground state of $\hat{H}$, and the expectation value $\langle\hat{O}\rangle_{\tau}$ reduces to the ground-state expectation value of $\hat{O}$.
Several classical approaches have been developed to approximate imaginary time evolution, most notably quantum Monte Carlo [RevModPhys.73.33] and tensor-network algorithms [Schollw_ck_2011]. In recent years, quantum algorithms for approximating imaginary time evolution on quantum hardware have also been proposed. Approaches such as Quantum Imaginary Time Evolution (QITE) [Motta_2019] and its variational variant (VQITE) [McArdle_2019] implement imaginary time evolution directly on a quantum computer in the Schrödinger picture. For a pedagogical introduction to these methods, see Ref. [angléscastillo2025understandingquantumimaginarytime].
@@ II.2 Pauli Propagation
A challenge of Pauli Propagation is the exponential growth of terms in $O(t)$, as each rotation can generate new Pauli strings. To maintain the computational cost tractable, truncation strategies are employed:
@@
II.3 Imaginary Time Evolution via Pauli Propagation
Following a similar approach to real-time Pauli Propagation, we define a set of rules to evolve a Pauli operator under an imaginary time Pauli operator
$V_{Q}(\tau)=\exp\left(-\frac{\tau}{2}Q\right)=\cosh\!\left(\frac{\tau}{2}\right)\mathbb{I}-\sin\!\left(\frac{\tau}{2}\right)Q,$ (10)
@@ II.3 Imaginary Time Evolution via Pauli Propagation
The evolution of a Pauli operator $P$ under this imaginary time operator satisfies
@@
Comparison of ITPP, Trotterized ITE, and Exact ITE
Figure 1: Comparison of Pauli Propagation imaginary time evolution (ITPP) with Trotterized imaginary time evolution and exact imaginary time evolution for the one-dimensional transverse-field Ising model (TFIM) with $N=10$ spins, $J=1$, and $h=0.5$, using a Trotter step size $\Delta t=0.04$. The left panel shows the evolution of the energy as a function of imaginary time, while the right panel displays the corresponding relative error with respect to the exact imaginary time evolution, shown on a logarithmic scale. The comparison illustrates the accuracy of ITPP relative to standard Trotterized evolution and the exact result over the course of imaginary time propagation.
where $\cdot\,$$\cdot\,$] and $\{\,\cdot\,$$\cdot\,\}$ denote the commutator and anticommutator, respectively. The derivation of these propagation rules is provided in Appendix A.
Interestingly, the propagation rules for a Pauli string under imaginary time evolution are essentially the opposite of those encountered in real-time Pauli Propagation. In particular, when the propagated Pauli string anticommutes with the generator of the imaginary time operator, the operator remains unchanged. By contrast, when the two operators commute, the Pauli string splits into two terms.
This behavior contrasts with the real-time case, where nontrivial evolution arises from anticommuting generators and the resulting coefficients are trigonometric functions reflecting unitarity and norm preservation. In imaginary time evolution, the coefficients are instead hyperbolic functions, corresponding to the non-unitary nature of the dynamics.
Together, these rules provide a natural extension of Pauli Propagation to imaginary time evolution, enabling the simulation of non-unitary dynamics and the estimation of thermal and ground-state properties within the Pauli operator framework.
Similar to real-time evolution, consider a Hamiltonian decomposed as $\hat{H}=\sum_{j}\alpha_{j}h_{j}$, where $h_{j}$ are Pauli operators. The exact imaginary time evolution operator,
$V(\tau)=e^{-\tau\hat{H}}$, (12)
can be approximated using a first-order Trotter expansion. Writing $\tau=n\,\Delta\tau$, we define the Trotterized evolution operator
$V_{\mathrm{Trotter}}(\tau)=\left(\prod_{j}\exp\!\left(-\alpha_{j}\Delta\tau\,h_{j}\right)\right)^{n}.$ (13)
@@
This decomposition expresses the imaginary time propagator as a sequence of single-Pauli imaginary time operators, each applied iteratively to propagate a Pauli operator. Combined with the propagation rules described above, this approximation enables the implementation of imaginary time evolution within the Pauli propagation framework. In particular, the thermal state of Eq. (4) can be approximated by propagating the maximally mixed state $\frac{1}{2^{n}}\,\mathbb{I}$ through the Trotterized imaginary time propagator, implemented as a sequence of symmetric half-step imaginary time evolutions generated by the individual Pauli terms of the Hamiltonian, with proper normalization. Applying Pauli Propagation in this setting and normalizing by the trace of the resulting operator yields an approximate thermal state
@@
$\tilde{\rho}_{\mathrm{th}}(\tau)=\frac{\text{PP}\!\left[V_{Trotter}(\frac{\tau}{2})\,\hat{I}\,V_{Trotter}(\frac{\tau}{2})\right]}{\operatorname{Tr}\!\left(\text{PP}\!\left[V_{Trotter}(\frac{\tau}{2})\,\hat{I}\,V_{Trotter}(\frac{\tau}{2})\right]\right)}.$ (15)
Here, $V_{\mathrm{Trotter}}(\tfrac{\tau}{2})$ denotes the Trotter approximation defined in Eq. (14), and $\text{PP}[\cdot]$ denotes the operator obtained after propagation and truncation in the Pauli basis. To control computational cost, a truncation scheme is applied during propagation, analogous to that employed in Pauli Propagation for real-time dynamics. In practice, the operator is normalized at each propagation step to ensure numerical stability and prevent uncontrolled growth of the Pauli coefficients.
Once the approximate state $\tilde{\rho}_{th}(\tau)$ is obtained, expectation values of observables $\hat{O}$ can be computed directly as
$\langle\hat{O}\rangle_{\tau}\approx\operatorname{Tr}\!\left[\hat{O}\,\tilde{\rho}_{th}(\tau)\right].$ (16)
This procedure enables the prediction of thermal or ground-state properties while retaining the flexibility to evaluate multiple observables using the same propagated operator. By adjusting the truncation threshold and the Trotter step size, one can systematically tune the tradeoff between computational cost and accuracy, making this approach well suited for controlled and flexible studies of quantum many-body systems.
It is important to note that truncation in the Pauli basis does not, in general, preserve positive semidefiniteness. Consequently, the approximate operator $\tilde{\rho}_{th}(\tau)$ may develop negative eigenvalues and therefore fail to represent a valid quantum state, rendering expectation values physically inconsistent. A straightforward way to restore positive semidefiniteness is to square the approximate density operator. This guarantees a positive semidefinite operator but comes at a significant computational cost, as squaring substantially increases the Pauli support. The resulting operator must then be normalized by its trace to recover a properly normalized quantum state. Since the procedure effectively applies the imaginary time evolution twice, the squared operator corresponds to an approximate thermal state at inverse temperature $2\tau$.
The computational overhead associated with explicitly enforcing positive semidefiniteness can be mitigated using the following identity. Here, $\langle\hat{O}\rangle_{2\tau}$ denotes the expectation value of the observable $\hat{O}$ at imaginary time $2\tau$, and $\tilde{\rho}_{\mathrm{th}}(\tau)$ is the ITPP approximation to the thermal state at imaginary time $\tau$:
$\langle\hat{O}\rangle_{2\tau}\approx\frac{\operatorname{Tr}\!\left[\hat{O}\,\tilde{\rho}_{\mathrm{th}}^{2}(\tau)\right]}{\operatorname{Tr}\!\left[\tilde{\rho}_{\mathrm{th}}^{2}(\tau)\right]}=\frac{\operatorname{Tr}\!\left[\big(\hat{O}\,\tilde{\rho}_{\mathrm{th}}(\tau)\big)\tilde{\rho}_{\mathrm{th}}(\tau)\right]}{\operatorname{Tr}\!\left[\tilde{\rho}_{\mathrm{th}}^{2}(\tau)\right]}.$ (17)
Rather than explicitly constructing $\tilde{\rho}_{\mathrm{th}}^{2}(\tau)$, one can instead compute the Pauli expansion of the product $\hat{O}\,\tilde{\rho}_{\mathrm{th}}(\tau)$—which typically contains significantly fewer terms—and subsequently evaluate its overlap with $\tilde{\rho}_{\mathrm{th}}(\tau)$. The denominator corresponds to the purity of the approximate state and can be computed efficiently from the $L_{2}$ norm of the vector in the Pauli basis. This procedure yields physically meaningful approximations to the expectation values of the observables of interest.
When evolved for sufficiently long imaginary time, this framework enables the extraction of ground-state properties of observables of interest.
ITPP performance for $N=12$ in the ordered phase ($J=1$, $h=0.5$)
@@
Figure 2: Performance of ITPP with coefficient-threshold truncation for various thresholds in the ordered-phase TFIM ($J=1$, $h=0.5$) with $N=12$ spins and Trotter step $\Delta t=0.04$. The three panels show, from left to right, the energy during imaginary time evolution, the relative error with respect to the exact ground-state energy (displayed on a logarithmic scale), and the number of Pauli terms retained under different truncation thresholds. In the left-most panel, the solid black line indicates the exact ground-state energy computed using the Bogoliubov–de Gennes solution. The results illustrate the trade-off between accuracy and computational cost controlled by the truncation threshold.

### v2

# Pauli Propagation for Imaginary-Time Evolution

@@ Abstract
We extend the Pauli Propagation framework to simulate imaginary-time evolution. By deriving explicit update rules for the propagation of Pauli operators under imaginary-time evolution generated by Pauli strings, we introduce an imaginary-time Pauli Propagation (ITPP) algorithm for approximating imaginary-time dynamics directly in the Pauli basis. This approach enables the computation of thermal and ground-state properties while retaining the key computational advantages of Pauli Propagation. Benchmarking ITPP on the one-dimensional transverse-field Ising model demonstrates that truncation provides a controlled trade-off between accuracy and computational cost, while also revealing challenges associated with operator growth under imaginary-time evolution. Finally, combining imaginary-time and real-time Pauli Propagation naturally suggests a pathway toward simulating open quantum system dynamics within a unified framework.
@@ I Introduction
Classical simulation of quantum many-body systems is notoriously challenging due to the exponential growth of the Hilbert-space dimension with system size, rendering exact diagonalization rapidly intractable. As a consequence, approximate numerical methods are indispensable. Established techniques such as tensor network algorithms [Schollw_ck_2011, M_ller_Hermes_2012, Or_s_2014, Ba_uls_2023] and quantum Monte Carlo [RevModPhys.73.33, Assaad2008] have enabled substantial progress, yet they also face well-known limitations, including entanglement growth in out-of-equilibrium dynamics [RevModPhys.82.277, Calabrese_2005] and sign-problem constraints in frustrated or fermionic models [PhysRevB.41.9301, PhysRevLett.94.170201].
@@
The original formulation of Pauli Propagation applies to real-time dynamics. In this work, we extend Pauli Propagation to imaginary-time evolution. imaginary-time evolution (ITE) is a fundamental tool for state preparation, enabling access to thermal states and filtering toward ground states, and is routinely implemented in classical simulation methods such as quantum Monte Carlo [RevModPhys.73.33] and tensor-network algorithms [Schollw_ck_2011]. We derive explicit rules governing the evolution of Pauli operators under imaginary-time evolution generated by arbitrary Pauli strings. Building on these rules, we construct an algorithm for approximating thermal and ground-state observables. As in real-time Pauli Propagation, the number of Pauli terms typically grows rapidly under imaginary-time evolution, requiring truncation strategies to retain computational feasibility. Moreover, the ability to handle non-unitary imaginary-time evolution within the Pauli Propagation framework opens the door to combining imaginary- and real-time dynamics, thus enabling the simulation of open quantum system dynamics [PRXQuantum.3.010320].
To benchmark the method and to assess its practical performance, we apply the imaginary-time evolution Pauli Propagation (ITPP) algorithm to the one-dimensional transverse-field Ising model in the ordered phase. We compare its results with Trotterized imaginary-time evolution and exact calculations, and we examine several truncation strategies, including coefficient thresholds and fixed-rank schemes.
The paper is organized as follows. In Sec. II, we briefly review imaginary-time evolution, thermal states, and the standard Pauli Propagation algorithm for real-time dynamics. Next, we introduce our extension of Pauli Propagation to imaginary-time dynamics. Numerical experiments are presented in Sec. III. Finally, Sec. IV summarizes our findings and discusses possible future directions.
@@ II Theory
In this section, we review the theoretical background underlying our approach. We begin by summarizing imaginary-time evolution and its connection to thermal and ground states. We then outline the standard Pauli Propagation algorithm for real-time dynamics. Finally, we introduce our extension of Pauli Propagation to imaginary-time evolution, which forms the basis of the method used in this work.
II.1 Imaginary-Time Evolution
@@
More generally, imaginary-time evolution can be formulated at the level of density matrices. Given an initial density matrix $\rho_{0}$, its imaginary-time evolved state is defined as
$\rho(\tau)=\frac{e^{-\tau H}\,\rho_{0}\,e^{-\tau H}}{\operatorname{Tr}\!\left[e^{-2\tau H}\rho_{0}\right]}.$ (2)
This expression ensures proper normalization, $\operatorname{Tr}[\rho(\tau)]=1$, at all imaginary-times.
@@
In this case, the imaginary-time evolution defined in Eq. (2) yields, for an arbitrary imaginary-time $\tau$, the thermal state
$\rho_{\mathrm{th}}(\tau)=\frac{e^{-2\tau H}}{\operatorname{Tr}\!\left[e^{-2\tau H}\right]},$ (4)
obtained by evolving the infinite-temperature state at $\tau=0$ to imaginary-time $\tau$, which sets the inverse temperature $\beta=2\tau$. As discussed above, in the zero-temperature limit, $\tau\to\infty$, the imaginary-time evolution converges to a projector onto the ground state.
Expectation values of observables $O$ at imaginary-time $\tau$ are obtained as
$\langle O\rangle_{\tau}=\operatorname{Tr}\!\left[O\,\rho_{th}(\tau)\right]=\frac{\operatorname{Tr}\!\left[O\,e^{-2\tau H}\right]}{\operatorname{Tr}\!\left[e^{-2\tau H}\right]}.$ (5)
In the zero-temperature limit $\tau\to\infty$, the thermal state $\rho_{th}(\tau)$ projects onto the ground state of $H$, and the expectation value $\langle O\rangle_{\tau}$ reduces to the ground-state expectation value of $O$.
Several classical approaches have been developed to approximate imaginary-time evolution, most notably quantum Monte Carlo [RevModPhys.73.33] and tensor-network algorithms [Schollw_ck_2011]. In recent years, quantum algorithms for approximating imaginary-time evolution on quantum hardware have also been proposed. Approaches such as Quantum imaginary-time Evolution (QITE) [Motta_2019] and its variational variant (VQITE) [McArdle_2019] implement imaginary-time evolution directly on a quantum computer in the Schrödinger picture. For a pedagogical introduction to these methods, see Ref. [angléscastillo2025understandingquantumimaginarytime].
@@ II.2 Pauli Propagation
A challenge of Pauli Propagation is the exponential growth of terms in $O(t)$, as each rotation can generate new Pauli strings. To maintain the computational cost tractable, two main truncation strategies are employed:
@@
II.3 Imaginary-Time Evolution via Pauli Propagation
Following a similar approach to real-time Pauli Propagation, we define a set of rules to evolve a Pauli operator under an imaginary-time Pauli operator
$V_{Q}(\tau)=\exp\left(-\frac{\tau}{2}Q\right)=\cosh\!\left(\frac{\tau}{2}\right)\mathbb{I}-\sinh\!\left(\frac{\tau}{2}\right)Q,$ (10)
@@ II.3 Imaginary-Time Evolution via Pauli Propagation
The evolution of a Pauli operator $P$ under this imaginary-time operator satisfies
@@
The derivation of these propagation rules is provided in Appendix A.
Interestingly, the propagation rules for a Pauli string under imaginary-time evolution are essentially the opposite of those encountered in real-time Pauli Propagation. In particular, when the propagated Pauli string anticommutes with the generator of the imaginary-time operator, the operator remains unchanged. By contrast, when the two operators commute, the Pauli string splits into two terms.
This behavior contrasts with the real-time case, where nontrivial evolution arises from anticommuting generators and the resulting coefficients are trigonometric functions reflecting unitarity and norm preservation. In imaginary-time evolution, the coefficients are instead hyperbolic functions, corresponding to the non-unitary nature of the dynamics.
Together, these rules provide a natural extension of Pauli Propagation to imaginary-time evolution, enabling the estimation of thermal and ground-state properties within the Pauli operator framework.
Similar to real-time evolution, consider a Hamiltonian decomposed as $H=\sum_{j}\alpha_{j}h_{j}$, where $h_{j}$ are Pauli operators. The exact imaginary-time evolution operator,
$V(\tau)=e^{-\frac{\tau}{2}H},$ (12)
can be approximated using a first-order Trotter expansion. In defining the imaginary-time evolution operator, we include the factor of $1/2$ in the exponent so that, when the operator is applied on both sides of the maximally mixed initial state as $V(\tau)\,\frac{\mathbb{I}}{2^{n}}\,V^{\dagger}(\tau)$, the imaginary-time parameter $\tau$ directly corresponds to the inverse temperature $\beta$ that defines the thermal state. Writing $\tau=n\,\Delta\tau$, we define the Trotterized evolution operator
$V_{\mathrm{Trotter}}(\tau)=\left(\prod_{j}\exp\!\left(-\frac{\alpha_{j}}{2}\Delta\tau\,h_{j}\right)\right)^{n}.$ (13)
@@
This decomposition expresses the imaginary-time propagator as a product of single-Pauli imaginary-time evolution operators. Consequently, imaginary-time evolution can be implemented by successively propagating operators expressed in the Pauli basis using the propagation rules introduced in Eq. (11).
In order to compute thermal and ground-state properties, we approximate the thermal state of Eq. (4) by propagating the maximally mixed state $\frac{1}{2^{n}}\mathbb{I}$ under the Trotterized imaginary-time evolution operator,
@@
$\rho_{\mathrm{th}}(\frac{\tau}{2})\approx\frac{V_{\mathrm{Trotter}}^{\dagger}(\tau)\,\frac{\mathbb{I}}{2^{n}}\,V_{\mathrm{Trotter}}(\tau)}{\operatorname{Tr}\!\left[V_{\mathrm{Trotter}}^{\dagger}(\tau)\,\frac{\mathbb{I}}{2^{n}}\,V_{\mathrm{Trotter}}(\tau)\right]}.$ (15)
In practice, we can omit the preceding factor $\frac{1}{2^{n}}$ of the identity operator, as it appears both in the numerator and denominator. So, we can propagate directly the identity operator, yielding the same results.
This propagation can be carried out using the rules introduced in Eq. (11). In the first step, the identity operator is propagated through the initial imaginary-time operator. Since the identity commutes with any Pauli string, this operation preserves the identity while generating an additional Pauli string associated with the generator of the imaginary-time evolution. For generality, we therefore express the resulting operator as
$\rho=\sum_{\alpha}c_{\alpha}P_{\alpha},$
a linear combination of Pauli strings $P_{\alpha}$.
We then propagate $\rho$ under the next imaginary-time step generated by a Pauli string $Q$,
$V_{Q}(\tau)=\exp\!\left(-\frac{\tau}{2}Q\right),$ (16)
leading to the transformation
$\rho\;\longrightarrow\;V_{Q}^{\dagger}(\tau)\,\rho\,V_{Q}(\tau).$ (17)
Since $\rho$ is a sum of Pauli strings, the propagation rules (Eq. (11)) can be applied independently to each term $P_{\alpha}$. This produces an updated operator that is again a linear combination of Pauli strings. The operator is then re-expanded in the Pauli basis and a truncation scheme is applied to limit the growth of the number of Pauli terms, thereby controlling computational cost. In addition, the operator is normalized by its trace after each propagation step to ensure numerical stability and to prevent uncontrolled amplification of coefficients during imaginary-time evolution.
The resulting operator serves as the input for the next imaginary-time operation in the Trotter decomposition. Repeating this procedure for all Pauli terms in each Trotter step, and iterating over the full sequence of Trotter steps, defines the ITPP algorithm and yields an approximate implementation of imaginary-time evolution within the Pauli propagation framework. A step-by-step description of ITPP is provided in Algorithm II.3.
{algorithm}
[t] Imaginary-Time Pauli Propagation (ITPP)
1:Input: Hamiltonian $H=\sum_{k}h_{k}Q_{k}$; step size $\Delta\tau$; final imaginary-time $\tau$
2:Number of Trotter steps: $N_{\tau}=\left\lceil\tau/\Delta\tau\right\rceil$
3:Trotterize imaginary-time evolution:
$e^{-\frac{\tau}{2}H}\approx\left(\prod_{k}e^{-\frac{\Delta\tau}{2}h_{k}Q_{k}}\right)^{N_{\tau}}$
4:Initialize $\rho=\mathbb{I}$.
5:for Trotter step $n=1,\dots,N_{\tau}$ do
6:for each Pauli term $h_{k}Q_{k}$ in the chosen Trotter ordering do
7:  $\rho^{\prime}\leftarrow 0$
8:  for each Pauli term $c_{\alpha}P_{\alpha}$ in $\rho$ do
9:   Propagate using ITPP rules:
$P_{\alpha}\rightarrow V_{Q_{k}}^{\dagger}P_{\alpha}V_{Q_{k}}$
10:   Accumulate resulting Pauli strings into $\rho^{\prime}$
11:  end for
12:  Truncate and normalize $\rho\leftarrow\mathcal{T}(\rho^{\prime})/\mathrm{Tr}(\mathcal{T}(\rho^{\prime}))$
13:end for
14:end for
15:Output: propagated operator $\rho$
Comparison of ITPP, Trotterized ITE, and Exact ITE
@@
Figure 1: Comparison of ITPP with Trotterized and exact imaginary-time evolution for the one-dimensional transverse-field Ising model (TFIM) with $N=10$ spins, $J=1$, and $h=0.5$. The simulations use a Trotter step size $\Delta\tau=0.04$ and truncation threshold $\delta=0$. The left panel shows the energy as a function of imaginary time, while the right panel displays the corresponding relative error with respect to the exact ground-state energy computed using the Bogoliubov–de Gennes solution. In the left panel, the solid line indicates this exact ground-state energy. The relative error is shown on a logarithmic scale. The results demonstrate that ITPP faithfully reproduces Trotterized imaginary-time evolution.
Once the approximate state $\tilde{\rho}_{th}(\tau)$ is obtained, expectation values of observables $O$ can be computed directly as
$\langle O\rangle_{\tau}\approx\operatorname{Tr}\!\left[O\,\tilde{\rho}_{th}(\tau)\right].$ (18)
This procedure enables the prediction of thermal or ground-state properties, while retaining the flexibility to evaluate multiple observables using the same propagated operator. By adjusting the truncation threshold and the Trotter step size, one can systematically tune the tradeoff between computational cost and accuracy, making this approach well suited for controlled and flexible studies of quantum many-body systems.
It is important to note that truncation in the Pauli basis does not, in general, preserve positive semidefiniteness. Consequently, the approximate operator $\tilde{\rho}_{th}(\tau)$ may develop negative eigenvalues and therefore fail to represent a valid quantum state, rendering expectation values physically inconsistent. A straightforward way to restore positive semidefiniteness is to square the approximate density operator. This guarantees a positive semidefinite operator but comes at a significant computational cost, as squaring substantially increases the Pauli support. The resulting operator must then be normalized by its trace to recover a properly normalized quantum state. Since the procedure effectively applies the imaginary-time evolution twice, the squared operator corresponds to an approximate thermal state at inverse temperature $\beta=2\tau$.
The computational overhead associated with explicitly enforcing positive semidefiniteness can be mitigated using the following identity. Here, $\langle O\rangle_{2\tau}$ denotes the expectation value of the observable $O$ at imaginary-time $2\tau$, and $\tilde{\rho}_{\mathrm{th}}(\tau)$ is the ITPP approximation to the thermal state at imaginary-time $\tau$:
$\langle O\rangle_{2\tau}\approx\frac{\operatorname{Tr}\!\left[O\,\tilde{\rho}_{\mathrm{th}}^{2}(\tau)\right]}{\operatorname{Tr}\!\left[\tilde{\rho}_{\mathrm{th}}^{2}(\tau)\right]}=\frac{\operatorname{Tr}\!\left[\big(O\,\tilde{\rho}_{\mathrm{th}}(\tau)\big)\tilde{\rho}_{\mathrm{th}}(\tau)\right]}{\operatorname{Tr}\!\left[\tilde{\rho}_{\mathrm{th}}^{2}(\tau)\right]}.$ (19)
Rather than explicitly constructing $\tilde{\rho}_{\mathrm{th}}^{2}(\tau)$, one can instead compute the Pauli expansion of the product $O\,\tilde{\rho}_{\mathrm{th}}(\tau)$—which typically contains significantly fewer terms—and subsequently evaluate its overlap with $\tilde{\rho}_{\mathrm{th}}(\tau)$. The denominator corresponds to the purity of the approximate state and can be computed efficiently from the $L_{2}$ norm of the vector in the Pauli basis. This procedure yields physically meaningful approximations to the expectation values of the observables of interest.